How do I solve this problem?
1 1/15 + 3 3/10 - 2 4/5 =

Answer by checkley71(8403)

1 1/15+3 3/10-2 4/5 = 16/15+33/10-14/5 OR USING A COMMON DENOMINATOR OF 30 WE GET 32/30+99/30-84/30=(32+99-84)/30 OR 47/30 OR 1 17/30